OpenStudy (anonymous):

Dana was picking strawberries to sell and placing them in baskets on a table. The number of strawberries in the basket is 4 more than the number of baskets on a table. If b represents the number of baskets on each table, which expression can be used to represent the number of strawberries on 5 tables? 5b4 5(b + 4) 5b(b + 4) 5b + 4

OpenStudy (cwrw238):

b is number of baskets on 1 table so how many baskets are there on 5 tables? also number of strawberries in 1 basket = (b + 4)

OpenStudy (cwrw238):

how do you write 5 times b in algebra? thats the total number of baskets

OpenStudy (anonymous):

i dont know but i got it thanks :)

OpenStudy (anonymous):

try this one Evaluate the following expression using the values given: Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3.

OpenStudy (cwrw238):

???

OpenStudy (cwrw238):

all you do to evaluate the second one is to replace x by 3, y by 5 and z by -3 and work it out so its (3)^3 - 2(5)^2 - 3(3)^3 + (-3)^4

OpenStudy (cwrw238):

you work out powers this way (2)^4 = 2*2*2*2 = 16 (-2)^2 = -2 * -2 = 4

OpenStudy (cwrw238):

x3 = x^3 =(3)^3 = 3*3*3 = 27 that's the first part done for you

OpenStudy (anonymous):

its one of these -185 -77 -23 85

OpenStudy (cwrw238):

ok so what i would like you to do is work out the rest of the terms and see what you get I could just tell you the answer but if i do that you won't learn anything

OpenStudy (cwrw238):

-2(5)^2 = -2*5*5 = ?

OpenStudy (anonymous):

i got -50

OpenStudy (cwrw238):

yes thats right so we have 27 - 50 - 3(27) + (-4)^4 -3*27 = -81 and + 4(-4)^4 = + 256 so its 27 - 50 - 81 + 256

OpenStudy (cwrw238):

oh sorry its (-3^4 for the last one so 27 - 50 - 81 +(-3)4 -3*-3*-3*-3 = 81 so its 27 - 50 - 81+ 81

OpenStudy (cwrw238):

ok can you finish it off? i'll start it for you 27 - 50 -81 + 81 = -23 -81+81 = ?

OpenStudy (cwrw238):

sorry i was so long answering - my notifications are not updating